Let a 4x4 matrix M =
7 5 -1 8 //
7 5 -2 3 //
7 5 5 9 //
7 5 6 10

Find a 3 x 3 matrix A of the form
3 -1 -5
3 * *
3 * *

such that det(M) = 7 det(A)
1
Expert's answer
2020-03-19T17:06:22-0400

If we carefuly look through the first matrix we can easily understand that the det(M) = 0. It folows from the properties of the determinant (multiplying the second column by (7/5) and adding to the first column, we get the first column of zeros. It means the determinant is 0).

So that exactly means we need for second matrix determinant = 0 (det(M) = 0 = 7 det(A) "\implies"det(A) = 0) . So there many ways of the solution but it does not say to find all of them. For several of them algorithm is next - from the property that we've already used erlier determinant = 0, when one of the column is combination of the others. So we can put -1 into places (2,2) and (3,2) and we don't care about the others values or we can also put -5 into places (2,3) and (3,3) and as well don't care about the rest.

The answer is "\begin{vmatrix}
 3 & -1 & -5\\
 3 & -1 & a1 \\
 3 & -1 & a2
\end{vmatrix}" or "\begin{vmatrix}
 3 & -1 & -5\\
 3 & a1 & -5 \\
 3 & a2 & -5
\end{vmatrix}" where a1, a2 - parametrs (any number can be used).

As example "\begin{vmatrix}
 3 & -1 & -5\\
 3 & -1 & 1 \\
 3 & -1 & 2
\end{vmatrix}"